OpenStudy (anonymous):

Can someone help i have no clue how to do this.. What is the graph of the function f(x) = the quantity of x plus 7, all over x minus 4?

OpenStudy (mathmale):

If at all possible, please present your equations symbolically, not in words. Your "function f(x) = the quantity of x plus 7, all over x minus 4" should look like:\[f(x)=\frac{ x+7 }{ x-4 }\]

OpenStudy (anonymous):

first time on here didnt know

OpenStudy (mathmale):

Check this out: what happens to f(x) if x=4? It's important that you recognize this. Secondly, what happens as x grows larger and larger? y approaches what value? Try x=10, x=100, x=1000, and so on, and keep track of the resulting y-values.

OpenStudy (mathmale):

I'm glad you're on OpenStudy now and hope you will find it to be a great place to learn and teach.

OpenStudy (anonymous):

im still confused

OpenStudy (mathmale):

Have you heard of or used the word "asymptote" before?

OpenStudy (mathmale):

It has to do with graphing functions.

OpenStudy (mathmale):

I asked you to try substituting x=4 into the given equation and tell me what happens to the denominator. What happens to the whole fraction?

OpenStudy (anonymous):

it cancels out?

OpenStudy (mathmale):

Hope you continue to find it interesting to explore OpenStudy. To be honest with you, I'm expecting guests to arrive soon and will then be unavailable. I will help you all I can; when not here, someone else will likely step in.

OpenStudy (mathmale):

What cancels out? Explain.

OpenStudy (mathmale):

Elisha?

OpenStudy (anonymous):

the denominator would be zero

OpenStudy (anonymous):

4-4 right?

OpenStudy (mathmale):

Yes, and what happens if the den. becomes zero?

OpenStudy (mathmale):

11/0 = ?

OpenStudy (anonymous):

11?

OpenStudy (mathmale):

Is division by zero permitted?

OpenStudy (mathmale):

11/1=11; 11/0= ? 11/2 = ?

OpenStudy (anonymous):

no so would it be 0?

OpenStudy (mathmale):

Division by zero is "undefined" (not zero).

OpenStudy (mathmale):

the reason I'm leading you thru all this is that we cannot let x=4 in your function, because division by zero would occur. x may be less than 4 or more than 4, but not = to 4. Can you use the Draw utility to sketch a set of coordinate axes and then draw a vertical dashed line thru x=4?

OpenStudy (anonymous):

im not good with graphs at all

OpenStudy (mathmale):

Click on the Draw button, below. experiment a bit (use the menu). Then draw a set of coordinate axes and mark x=4 and draw a vert. line thru x=4. Practice is the way to become "good with graphs." Seriously.

OpenStudy (anonymous):

can you give me an idea?

OpenStudy (mathmale):

sorry for the delay. I have visitors fromj England here at my house in California. I will start you out with a simple drawing: